OpenStudy (anonymous):

Convert 0.0075 to a percent. A) 0.075% B) 0.75% C) 0.0075% D) 75%

OpenStudy (akashdeepdeb):

Percent should tell you that it is something in 1 hundreds. Thus 'a' percent = a/100 = 0.0a Thus to convert 0.0075 to a percent. x/100 = 0.0075 Thus x = 0.0075*100 Got it? :)